OpenStudy (anonymous):

find dy/dx for y=6^x-1

OpenStudy (luigi0210):

Is it \(\Large y=6^{x-1}\)

Elsa213 (elsa213):

y = x^-6 is also equal to y = 1 / x^6 So now we apply the quotient law of derivation: Which states if we have a function y(x) = f(x) / g(x) then the derivative of y(x) = y '(x) =[ f '(x)*g(x) - f(x)*g '(x) ] / ( g(x) )^2 So y '(x) = Dx[ 1 ]* x^6 - 1 * Dx[x^6] / ( x^6 )^2 = 0 * x^6 - 1*6x^5 / x^12 = 0 - 6x^5 / x^12 = -6x^5 / x^12 = -6 / x^7 Hence: y '(x) = -6 / x^7